# Neural Point-based Volumetric Avatar: Surface-guided Neural Points for Efficient and Photorealistic Volumetric Head Avatar

### v1

@@ 4.2 Ablation Studies
Effect of using different numbers of points. We investigate the impact of using different numbers of points on the rendering quality and inference time in Tab. 3 (a.1)-(a.3). The visual results are shown in Fig. 9. When the point number is small (i.e., NPVA-256), NPVA generates poor results (46.45 MSE) and contains holes due to insufficient point resolution. With increased point number (NPVA-1k), we notice obvious improvement (from 46.45 to 26.36 MSE) and do not see holes, indicating the point resolution is sufficient. Further increasing the point number (NPVA-2k) results in perceptually very similar results and slightly worse MSE (27.36 vs. 26.36).
@@
Importance of using an extra displacement map. Introducing an extra displacement map reduces MSE from 26.36 to 23.70 (Tab. 3 (a.2)). A visual comparison is provided in Fig. 9. We can also see using a displacement map is much more effective than using more points. This is because using a displacement map enables a more flexible arrangement of the neural points so that they can not only move on the surface (i.e., along the tangent plane) but also move along the normal direction, forming a thicker shell with increased capacity (see Fig. 4).
@@
Influence of the lightweight radiance decoding process. Using the proposed lightweight radiance decoding in Sec. 3.1.3 not only greatly reduces the inference time (3129 vs. 482 ms), but also improves the quality (23.70 vs 24.51 MSE). A visual comparison is provided in Fig. 5. A possible explanation is that using networks with too much capacity may cause overfitting and hinders the generalization on novel expressions for dynamic scene modeling tasks.
@@
Figure 6: Effect of different depth-guided shading point sampling strategies. There are two depth levels around the orange mark. With pixel-wise depth sampling [16], the model generates “mesh-like” artifacts. In contrast, our patch-wise depth sampling samples around both the front and back depth levels, leading to more realistic results.
@@
Different shading point sampling methods. We investigate the impact of different shading point sampling methods on a given ray. Tab. 3 (c.1, c.2) present results obtained using different sampling methods. Our patch-wise depth-guided sampling method reduces MSE from 24.92 to 23.70 compared to pixel-wise depth-guided (c.2) proposed in  [16]. Using “Pixel-Depth” easily causes inaccuracies, especially for the jaw region that contains two different depth levels, resulting in a mesh-like beard rendering (see Fig. 6). Training with a naive sampling strategy is slower and could not give a similar result (406.49 MSE) compared to the depth-guided sampling methods when using only 20 sample points per ray ($\sim$200 in NeRF).
@@
GEP training strategy. We compare our GEP ray sampling strategy (Sec. 3.2.2) with a naive alternative that always uses uniformly sampled rays. GEP achieves lower MSE (23.70 vs. 30.08) in Tab. 3 (d). The training loss curves and visual comparisons are shown in Fig. 8. The model trained with the naive strategy converges to a sub-optimal solution. Although achieving satisfactory renderings in smooth facial regions (e.g., skins), it struggles to handle challenging facial regions (e.g., eyes and mouth interior). In contrast, the model trained with our GEP strategy allocates more computing budget to these difficult regions and obtains more realistic facial renditions.
4.3 Analysis on Volumetric Methods
@@
We compare with single-frame NeRF fitting [27], which can be viewed as the upper limit of different volumetric avatars. The results are shown in Tab. 2 and Fig. 7. On single-frame fitting, our NPVA generates high-fidelity results comparable to NeRF while being 70$\sim$ 70$\times$ 70 × faster (524 vs. 38392 ms) during inference. What’s more, our NPVA can handle dynamic scenes (a 49-frame sequence) effectively with minor performance drop. In Fig. 7, we notice NPVA also generates visually better results (e.g., sharper and more realistic reflection effects) than NeRF, possibly due to the help of coarse geometry prior and the perceptual loss.
5 Conclusion & Discussion
@@
Limitation. We rely on coarse mesh tracking for modeling and optimization, which generally works well but does not account for very long hair or diverse hairstyles, such as those not present in tested female subjects. When relaxing the regularization on the displacement map for these cases, our method tends to produce blurry results for novel expressions (Fig. 7b).
Figure 7: Qualitative Comparison with NeRF and our failure case. (a) shows our qualitative comparison with NeRF. When trained on one single frame, our NPVA achieves results even visually better than NeRF (i.e., sharper and more realistic reflection effects), particularly in challenging facial regions (i.e., eyes and mouth interior) possibly due to the help of coarse geometry prior and the perceptual loss. (b) is one failure example, in which the regions far outside the rasterized mask contain obvious artifacts. This result is obtained by relaxing the constraint applied on the displacement map during training. Since no prior knowledge about the hair is provided in the coarse mesh, we replace our depth-guided shading point sampling strategy with the standard sampling strategy used in NeRF.
Figure 8: Comparisons on Different Training Strategies. The left part shows the photometric losses and validation MSE during training with different training strategies, while the right part shows qualitative rendering results. The model using our GEP training strategy converges to a lower photometric loss and achieves lower validation MSE, thereby producing more realistic facial reditions. As expected, the model using our GEP training strategy achieves significant reduction of validation MSE during the E-stage.
Figure 9: Ablation Study on Different Point Number and Displacements. When the number of points is low (i.e., $256\times 256$), the model produces unsatisfactory results, even with holes present. As the point number grows to $1024\times 1024$, the model obtains sufficient points and fills holes, generating improved visual results. Increasing the point number to $2048\times 2048$ results in perceptually very similar facial renditions. At this time, introducing our displacements can further improve the model performance, generating more realistic results, particularly in facial challenging regions (i.e., eyes and mouth interior).
Figure 10: Qualitative Comparisons with State-of-the-Art Methods.

### v3

@@ 4.2. Ablation Studies
We investigate the impact of using different numbers of points on the rendering quality and inference time in Tab. 2 (a.1)-(a.3). The visual results are shown in Fig. 9. When the point number is small (i.e., NPVA-256), NPVA generates poor results (46.45 MSE) and contains holes due to insufficient point resolution. With increased point number (NPVA-1k), we notice obvious improvement (from 46.45 to 26.36 MSE) and do not see holes, indicating the point resolution is sufficient. Further increasing the point number (NPVA-2k) results in perceptually very similar results and slightly worse MSE (27.36 vs. 26.36).
Importance of using an extra displacement map.
@@
Introducing an extra displacement map reduces MSE from 26.36 to 23.70 (Tab. 2 (a.2)). A visual comparison is provided in Fig. 9. We can also see using a displacement map is much more effective than using more points. This is because using a displacement map enables a more flexible arrangement of the neural points so that they can not only move on the surface (i.e., along the tangent plane) but also move along the normal direction, forming a thicker shell with increased capacity (see Fig. 4 and Fig. 10).
Influence of the lightweight radiance decoding process.
@@
Using the proposed lightweight radiance decoding in Sec. 3.1.3 not only greatly reduces the inference time (3129 vs. 482 ms), but also improves the quality (23.70 vs. 24.51 MSE). A visual comparison is provided in Fig. 5. A possible explanation is that using networks with too much capacity may cause overfitting and hinder the generalization of novel expressions for dynamic scene modeling tasks.
Figure 5. Effect of lightweight radiance decoding. Our lightweight radiance decoding not only gains $\sim 7\times$ speedup over Point-NeRF’s radiance decoding, but also produces better facial renditions for unseen expressions.
@@
Figure 6. Effect of different depth-guided shading point sampling strategies. There are two depth levels around the orange mark. With pixel-wise depth sampling (Lin et al., 2022), the model generates “mesh-like” artifacts. In contrast, our patch-wise depth sampling samples around both the front and back depth levels, leading to more realistic results.
Different shading point sampling methods.
@@
We investigate the impact of different shading point sampling methods on a given ray. Tab. 2 (c.1, c.2) present results obtained using different sampling methods. Our patch-wise depth-guided sampling method reduces MSE from 24.92 to 23.70 compared to pixel-wise depth-guided proposed in  (Lin et al., 2022). Using “Pixel-Depth” easily causes inaccuracies, especially for the jaw region with two different depth levels, resulting in a mesh-like beard rendering (see Fig. 6). Training with a naive sampling strategy is slower and could not give a similar result (406.49 MSE) compared to the depth-guided sampling methods when using only 20 sample points per ray ($\sim$200 in NeRF).
Different KNN search radius.
Increasing the KNN search radius from 3mm to 4mm improves rendering results (23.54 vs. 23.70 MSE) with longer processing time (563 vs. 482 ms). As shown in Fig. 8b, a larger search radius can eliminate holes in some extreme expressions.
GEP training strategy.
@@
We compare our GEP ray sampling strategy (Sec. 3.2.2) with a naive alternative that always uses uniformly sampled rays. GEP achieves lower MSE (23.70 vs. 30.08) in Tab. 2 (d). The training loss curves and visual comparisons are shown in Fig. 8a. The model trained with the naive strategy converges to a sub-optimal solution. Although achieving satisfactory renderings in smooth facial regions (e.g., skins), it struggles to handle challenging facial regions (e.g., eyes and mouth interior). In contrast, the model trained with our GEP strategy allocates more computing budget to these difficult regions and obtains more realistic facial renditions.
4.3. Analysis on Volumetric Methods
@@
We compare with single-frame NeRF fitting (Mildenhall et al., 2020), which can be viewed as the upper limit of different volumetric avatars. The results are shown in Tab. 3 and Fig. 7. On single-frame fitting, our NPVA generates high-fidelity results comparable to NeRF while being $\sim 70\times$ faster (524 vs. 38392 ms) during inference. What’s more, our NPVA can handle dynamic scenes (a 49-frame sequence) effectively with minor performance drops. In Fig. 7, we notice NPVA also generates visually better results (e.g., sharper and more realistic reflection effects) than NeRF, possibly due to the help of coarse geometry prior and the perceptual loss.
5. Conclusion & Discussion
@@
Limitation.
We rely on coarse mesh tracking for modeling and optimization, which generally works well but does not account for very long hair or diverse hairstyles, such as those not present in tested female subjects. When relaxing the regularization on the displacement map for these cases, our method tends to produce blurry results for novel expressions (Fig. 7b).
Acknowledgements.
This work was supported by the Natural Science Foundation of China (Project Number 62132012) and Tsinghua-Tencent Joint Laboratory for Internet Innovation Technology.